A car accelerates from 10.0 m/s to 30.0 m/s at a rate of 3.00 m/s2. How far does the car travel while accelerating?


a. 226 m

b. 133 m

c. 399 m

d. 80 m


1
Expert's answer
2022-02-15T16:14:53-0500

Given:

"v_1=10.0\:\rm m/s"

"v_2=30.0\:\rm m/s"

"a=3.00\:\rm m/s^2"


The distance traveled by a car

"d=\frac{v_2^2-v_1^2}{2a}=\rm \frac{30.0^2-10.0^2}{2*3.00}=133\:\rm m"

Answer: b. 133 m